Consider the perfectly competitive market for Diesel. The aggregate demand for gasoline is
The aggregate demand for gasoline is
Q_d=100-p
While the aggregate supply is Q_s= 3p

(i) It is established that most fuel stations are going out of business To address this problem, the government decides to set a minimum price of p ̅ = 30 . What will be the new equilibrium price and quantity? What will be the new consumer surplus and producer surplus? Who gains and who loses from this regulation? Explain how the total surplus affected? Briefly explain the intuition.
1
Expert's answer
2021-09-09T17:14:34-0400

At P=30

Q_d=100-30=70 units

3P=70

"P= \frac{70}{3}=23.33"

Consumer surplus "= \frac{1}{2} \times (70-0) \times (100-30)"

"= \frac{1}{2} \times 70 \times 70 \\

= 2450"

Producer surplus "= (70-0) \times (30-23.33)"

"= 70 \times 6.67 \\

= 466.9"